OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex (1,-108) and y-intercept(0,-105). PLEASE write anwser in the form of (x1,y1)(x2,y2) if it is necessary you can round to the nearest hundreth. <3 PLEASE HELP ME URGENT !

OpenStudy (anonymous):

ok lets do it

OpenStudy (anonymous):

ok :) yay !

OpenStudy (anonymous):

\[y=a(x-1)^2-108\] \[-105=a(0-1)^2-108\] \[-105=a-108\] \[-105+108=a\] \[a=3\] \[y=3(x-1)^2-108\]

OpenStudy (anonymous):

now set \[y=0\] solve for x

OpenStudy (anonymous):

\[0=3(x-1)^2-108\] \[108=3(x-1)^2\] \[36=(x-1)^2\] \[x-1=\pm\sqrt{36}\] \[x-1=\pm 6\] \[x-1=6\] \[x=7\] \[x-1=-6\] \[x=-5\]

OpenStudy (anonymous):

two solutions are \[(-5,0)\] and \[(7,0)\]

OpenStudy (anonymous):

ohkay so then after that we ...

OpenStudy (anonymous):

thanks <3

OpenStudy (anonymous):

yw